OpenStudy (anonymous):

A biologist finds and tags the 26 pigs she finds on an island. On her second visit, 10 of the 15 pigs she finds have tags. Which is the best estimate for the number of pigs on the island?

OpenStudy (campbell_st):

the method is called capute/recapture and used to estimate populations by capturing some animals, tagging and releasing the recapturing... It uses the idea of the recapture ratio of tagged will be consistent with the initial tagging process. so you need to use ratios \[\frac{population}{tagged}\] so \[\frac{15}{10} = \frac {x}{26}\] to estimate the population.. now just solve for x